# Holographic duality from Howe duality: Chern–Simons gravity as an ensemble of code CFTs

### v1

@@ 2.4.1 Relation to quantum stabilizer codes
The code CFT path integrals (53), as well as path integrals of the Chern-Simons theory on individual handlebody geometries, in terms of qudit Hilbert space ${\mathcal{H}}_{N}$ have a natural interpretation as quantum stabilizer states. This was first pointed out in Dymarsky:2024frx , [^12] The connection to CSS stabilizer states was independently made by A. Barbar. and is explained in more detail below. A relation between finite-field theta correspondence and quantum stabilizer codes was also noted in Montealegre-Mora2021 .
@@
Here ${\mathscr{C}}\subset{\mathbb{F}}_{p}^{2N}$ is an abelian group which is self-orthogonal (isotropic), meaning it is contained in its dual, ${\mathscr{C}}\subset{\mathscr{C}}^{\perp}$. The dual here is with respect to symplectic form, $[\rm x,\rm y]=0$ for any $\mathrm{x}$ and $\mathrm{y}\in{\mathscr{C}}$, where
@@
Comparing this with the genus one code CFT path integral states $\Psi_{\mathcal{C}}$ (39) we readily find that corresponding stabilizer group is defined by
@@
${\mathscr{C}}=\{(\eta a,b)\,|\,(a,b)=(\tilde{a},0)\gamma^{T}\,\,{\rm mod}\,\,p,\,\,{\tilde{a}}\in{\mathbb{F}}_{p}^{N}\},\quad f(\rm x)=0.$ (95)
@@
where ${\mathcal{L}}\subset{\mathbb{F}}_{p}^{2}$ is an additive classical symplectic self-dual code of length $22g=$ 2, such that for any $(a,b)\in{\mathcal{L}}$ and $(a^{\prime},b^{\prime})\in{\mathcal{L}}$,
$ab^{\prime}-a^{\prime}b\,\,{\rm mod}\,\,p=0,$ (97)
@@
${\mathscr{C}}=\{(\eta a_{1},\dots\eta a_{g},b_{1},\dots,b_{g})\,|\,(a_{1},\dots a_{g},b_{1},\dots,b_{g})=(\tilde{a}_{1}\dots\tilde{a}_{g},0,\dots 0,)\gamma^{T}\,\,{\rm mod}\,\,p,\,\,{\tilde{a}}_{i}\in{\mathscr{D}}\}.$ (102)
@@ 3 Howe duality and a proof of the averaging formula
Howe duality is a relation between representations of two Lie groups $G$ and $H$ over a field ${\mathbb{F}}$, embedded as commuting subgroups inside a larger symplectic group $\mathcal{S}$. In particular, this relation is a pairing or correspondence between the representations of $G$ and $H$ inside $G\times H\subset\mathcal{S}$, but considering different types of fields ${\mathbb{F}}$ leads to different precise statements. This correspondence was first introduced by Howe Howe1973Preprint ; Howe1979 ; Howe:1987tv ; Howe1989 ; Howe1989a based on earlier work by Siegel and Weil on modular forms Siegel1935 ; Siegel1951 ; Weil1964 ; Weil1965 and Segal and Shale on the quantization of the boson field Segal1959 ; Segal1961 ; Shale1962 .
@@
Here $\pi_{\lambda}$ are representations of $S_{g}$, and $\rho_{\lambda}$ are representations of $GL(n,{\mathbb{C}})$. They are both labeled by Young tableaux with $g$ boxes. For example, in the case of just two factors, we have $V\otimes V=S^{2}V+\Lambda^{2}V$, where $S^{2}V$ is the symmetric part (upon which $S_{2}$ acts trivially), and $\Lambda^{2}V$ is the antisymmetric part (upon which $S_{2}$ acts by a sign). Here representations $\pi_{\lambda}$ are the two one-dimensional irreps of $S_{2}$ and $\rho_{\lambda}$ are $n(n\pm 1)/2$ dimensional irreducible representations of $GL(n,{\mathbb{C}})$. [^13] Schur–Weyl duality has no direct relation to averages discussed in this work. However, see Maxfield:2023mdj for a setting where Schur–Weyl duality is central to characterizing the Hilbert space in a toy model of quantum gravity.
@@ 3.1 Generalities
Howe duality applies when $\mathcal{S}$ acts on a space $W$ and when its subgroups $G$ and $H$ form a dual pair, which requires that they be 1) reductive groups and 2) mutually centralizing inside $\mathcal{S}$. A group $G$ acts reductively on $W$ if it breaks $W$ into a finite sum of $G$-invariant subspaces. The centralizer of a subgroup $G\subset\mathcal{S}$ is the set of elements in $\mathcal{S}$ which commute with all $g\in G$. Howe gave a classification of reductive dual pairs Howe1979 ; Howe:1987tv ; Howe1989a :
@@
where $\mathbf{e}_{i}$ is a $2n\times g$ matrix with zeros except a single $1$ at a position determined by $1\leq i\leq 2gn$ (say row by row), and $\hat{X}_{i}$ is the operator $\hat{X}$ of (61) acting on the $i^{\text{th}}$ qudit. In the case of ${\mathbb{F}}={\mathbb{F}}_{p}$, each “creation operator” can only act $p$ times, and $\omega$ has dimension $p^{2ng}$. The oscillator representation $\omega$ is then isomorphic as a vector space to $\mathcal{X}$ introduced above.
The oscillator representation $\omega$ is an irreducible representation of the Heisenberg–Weyl group $H_{2ng}$. It also transforms projectively under the group of (outer) automorphisms of $H_{2ng}$, which is $Sp(4ng,F)$ – these are the canonical transformations of quantum mechanics. In case of field $F={\mathbb{R}}$, $\omega$ is a projective representation of $Sp(4ng,F)$ but a linear representation of the double cover $\widetilde{Sp}$ of the symplectic group, which is called the metaplectic group, and the pair $G\times H$ lifts to a pair of subgroups $\tilde{G}\times\tilde{H}$ of $\widetilde{Sp}$. [^14] For finite fields, it is possible to write down a genuine (non-projective) representation by including suitable phases, but this is impossible for ${\mathbb{R}}$. For the field ${\mathbb{C}}$, the oscillator is a genuine representation of $Sp(4gn,{\mathbb{C}})$ Howe1979 . A review of the details of this construction can be found in appendix A, where we give an introduction to the oscillator representation for both ${\mathbb{R}}$ and ${\mathbb{F}}_{p}$.
@@
$\omega|_{\tilde{G}\times\tilde{H}}=\sum_{i}\pi_{i}(\tilde{G})\otimes\rho_{i}(\tilde{H}).$ (109)
@@ 3.2 Finite-field Howe duality
where $\Theta(\pi)$ is not necessarily an irrep, meaning that the correspondence between irreps is not one-to-one. [^17] Some work has gone into understanding if there is a natural choice for a canonical representative irrep in $\Theta(\pi)$ which makes the correspondence one-to-one, see Gurevich2016 ; Aubert2016 ; Pan2020 . Since all involved spaces are finite, it is always possible to write the decomposition as (110) (if one allows for $\Theta(\pi_{i})=0$ for some $\pi_{i}$). The non-trivial part is therefore to characterize the map $\Theta$. A conjecture about $\Theta$ for a subclass of irreps, so-called unipotent representations, that include the singlet (trivial representation) was given in Aubert1996 and proven in Pan2019 ; Ma2022 . We give more details on this below, but first we formulate the main statement.
@@ 3.2.1 Proof of equality “Ensemble average equals Poincaré series”
$\phi_{\text{A}}=\frac{1}{\mathcal{N}}\sum_{C}\psi_{C},\qquad\phi_{\text{P}}=\frac{1}{\mathcal{N}^{\prime}}\sum_{h\in\Gamma\backslash Sp}U_{h}\psi_{0}$ (111)
It is easy to see that $\phi_{\text{A}}$ and $\phi_{\text{P}}$ are both invariant under both $H$ and $G$ [^18] The invariance of $\phi_{\text{A}}$ under $G$ and $\phi_{\text{P}}$ under $H$ are immediate since they are group averages. The invariance of $\phi_{\text{A}}$ under $H$ follows from modular invariance of each $\psi_{C}$. Finally, the invariance of $\phi_{\text{P}}$ under $G$ follows from commutativity of actions of $G$ and $H$ and the fact that $\psi_{0}$ is invariant under $G$. They are also both non-zero. [^19] Neither $\phi_{\text{A}}$ nor $\phi_{\text{P}}$ can vanish. Each term contributing to $\phi_{\text{A}}$ has a positive scalar product with $\psi_{0}$; each term contributing to $\phi_{\text{P}}$ has a positive scalar product with $\psi_{\mathcal{C}}$ for any code $\mathcal{C}$. Then we invoke the following lemma, proven below.
@@ 3.2.2 Theta map and proof of the lemma
We now introduce some definitions following Pan2020 , which will allow us to explain (114) from the statements of Aubert1996 ; Pan2019 ; Ma2022 . An important subclass of representations of both orthogonal and symplectic groups over ${\mathbb{F}}_{p}$ ($p$ odd) are so-called unipotent representations [^20] The definition of unipotent is complicated, see Definition 7.8 of Deligne1976 ; for us it suffices to know that this is a subclass of representations which includes the singlet of the respective groups. and can be labeled by “symbols.” This follows older results by Lusztig Lusztig1977 . A symbol is a pair of collections of strictly increasing numbers (or empty set),
@@
$\operatorname{rk}\Lambda=N,\qquad\operatorname{def}\Lambda=\begin{cases}1\text{ (mod $4$),}\quad&Sp(2N,{\mathbb{F}}_{p}),\\ 0\text{ (mod $4$),}\quad&O^{+}(2N,{\mathbb{F}}_{p}),\\ 2\text{ (mod $4$),}\quad&O^{-}(2N,{\mathbb{F}}_{p}).\\ \end{cases}$ (117)
@@
Unipotent representations map to unipotent representations under the $\Theta$ map. To give an explicit description of how, we change labels from symbols to bipartitions. A bipartition is a pair of Young tableaux, equivalently a pair of integer collections $[\kappa_{1},\ldots,\kappa_{m}]$, $\kappa_{i}\geqslant\kappa_{i+1}$. Restricting to the case of $\operatorname{def}\Lambda=1$ for $Sp$ and $\operatorname{def}\Lambda=0$ for $O^{+}$, we can map symbols to bipartitions by
@@
Let us now apply this consideration to the singlet (trivial) irrep. In terms of bi-partitions, the trivial irrep of $O^{+}(2n,{\mathbb{F}}_{p})$ has $(\kappa,\lambda)=([n],\emptyset)$, where $[n]$ denotes the Young tableau with a single row with $n$ boxes, see proposition 2.9 of Pan2020 . By checking the conditions in (121), we then find
@@ 3.3 Explicit example: $p=3$, $n=1$, $g=1$
Having discussed the general theory, we are ready to give an example with all the details spelled out. We will consider the simplest case, $p=3$, length $n=1$ and genus $g=1$. In appendix E, we give four additional examples: increasing each of the parameters (thus $[png]=[511],\,[321],\,[312]$), and finally the case $[png]=[322]$, where the sum (122) contains three terms instead of two as in the other examples.
@@
$W_{\mathcal{C}_{1}}=\psi_{00}+\psi_{10}+\psi_{20}\,,\qquad W_{\mathcal{C}_{2}}=\psi_{00}+\psi_{01}+\psi_{02}\,,\qquad$ (124)
@@
$SL_{2,3}$ $[1]$ $[4]$ $[4]^{\prime}$ $[1]^{\prime}$ $[4]^{\prime\prime}$ $[4]^{\prime\prime\prime}$ $[6]$
$r_{1}$ $1$ $1$ $1$ $1$ $1$ $1$ $1$
$r_{1}^{\prime}$ $1$ $\zeta^{2}$ $\zeta$ $1$ $\zeta$ $\zeta^{2}$ $1$
$\bar{r}_{1}^{\prime}$ $1$ $\zeta$ $\zeta^{2}$ $1$ $\zeta^{2}$ $\zeta$ $1$
$r_{2}$ $2$ $1$ $1$ $-2$ $-1$ $-1$ 00
$r_{2}^{\prime}$ $2$ $\zeta$ $\zeta^{2}$ $-2$ $-\zeta^{2}$ $-\zeta$ 00
$\bar{r}_{2}^{\prime}$ $2$ $\zeta^{2}$ $\zeta$ $-2$ $-\zeta$ $-\zeta^{2}$ 00
$r_{3}$ $3$ 00 00 $3$ 00 00 $-1$
$O_{2,3}^{+}$ $I$ $\left(\begin{smallmatrix}0&2\\ 2&0\end{smallmatrix}\right)$ $\left(\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right)$ $2I$
$\pi_{1}$ $1$ $1$ $1$ $1$
$\pi_{1}^{\prime}$ $1$ $-1$ $-1$ $1$
$\pi_{1}^{\prime\prime}$ $1$ $1$ $-1$ $-1$
$\pi_{1}^{\prime\prime\prime}$ $1$ $-1$ $1$ $-1$
@@
The larger symplectic group that acts on $\mathcal{X}$ is $\mathrm{Sp}(4,{\mathbb{F}}_{3})$. It has 51 8405184051\,84051 840 elements and 34 irreps. Four of these are “minimal” (see appendix B) of dimension $4,4,5,5$, the other non-trivial irreps have dimensions between $6$ and $81$. Our space $\mathcal{X}$, which is isomorphic to the oscillator representation, is here a direct sum of one irrep with dimension $5$ (charge-conjugation even) and one irrep with dimension $4$ (charge-conjugation odd) under $Sp(4,{\mathbb{F}}_{3})$.
@@
We finish this section with a comment on modular invariants. We see that the following two spaces
$\pi_{1}\otimes r_{1}$ $\leftrightarrow\operatorname{span}\{2\psi_{00}+\psi_{01}+\psi_{02}+\psi_{10}+\psi_{20}\}$ (134)
$\pi_{1}^{\prime}\otimes r_{1}$ $\leftrightarrow\operatorname{span}\{\psi_{01}+\psi_{02}-\psi_{10}-\psi_{20}\}$ (135)
@@
This idea that the enumerator polynomials of codes span all invariants of the MacWilliams identities (generally denoted Clifford groups) is an important property in the study of error-correcting codes Nebe2006book . However, it is known that codes cannot be used to generate all modular forms (at large enough genus), even in the chiral case, as discussed in SalvatiManni1986 ; Runge1996 ; Oura2008 .

### v2

@@ 2.4.1 Relation to quantum stabilizer codes
The code CFT path integrals (53), as well as path integrals of the Chern–Simons theory on individual handlebody geometries, in terms of qudit Hilbert space ${\mathcal{H}}_{N}$ have a natural interpretation as quantum stabilizer states. This was first pointed out in Dymarsky:2024frx , [^14] The connection to CSS stabilizer states was independently made by A. Barbar. and is explained in more detail below. A relation between finite-field theta correspondence and quantum stabilizer codes was also noted in Montealegre-Mora2021 .
@@
Here ${\mathscr{C}}\subset{\mathbb{F}}_{p}^{2N}$ is an abelian group which is self-orthogonal (isotropic), meaning it is contained in its dual, ${\mathscr{C}}\subset{\mathscr{C}}^{\perp}$. The dual here is with respect to symplectic form, $\rm x,\rm y]=0$ for any $\mathrm{x}$ and $\mathrm{y}\in{\mathscr{C}}$, where
@@
Comparing this with the genus-one code CFT path integral states $\Psi_{\mathcal{C}}$ (39) we readily find that corresponding stabilizer group is defined by
@@
${\mathscr{C}}=\{(\eta a,b)\,|\,(a,b)=(\tilde{a},0)\gamma^{T}\,\,({\rm mod}\,\,p),\,\,{\tilde{a}}\in{\mathbb{F}}_{p}^{N}\},\quad f(\rm x)=0.$ (95)
@@
where ${\mathcal{L}}\subset{\mathbb{F}}_{p}^{2}$ is an additive classical symplectic self-dual code of length $2g=2$, such that for any $(a,b)\in{\mathcal{L}}$ and $(a^{\prime},b^{\prime})\in{\mathcal{L}}$,
$ab^{\prime}-a^{\prime}b\,\,({\rm mod}\,\,p)=0,$ (97)
@@
${\mathscr{C}}=\{(\eta a_{1},...,\eta a_{g},b_{1},...,b_{g})\,|\,(a_{1},...,a_{g},b_{1},...,b_{g})\!=\!(\tilde{a}_{1},...,\tilde{a}_{g},0,...,0)\gamma^{T}\,\,({\rm mod}\,\,p),\,\,{\tilde{a}}_{i}\in{\mathscr{D}}\}.$ (102)
@@ 3 Howe duality and a proof of the averaging formula
Howe duality is a relation between representations of two Lie groups $G$ and $H$ over a field $F$, embedded as commuting subgroups inside a larger symplectic group $\mathcal{S}$. In particular, this relation is a pairing or correspondence between the representations of $G$ and $H$ inside $G\times H\subset\mathcal{S}$, but considering different types of fields $F$ leads to different precise statements. This correspondence was first introduced by Howe Howe1973Preprint ; Howe1979 ; Howe:1987tv ; Howe1989 ; Howe1989a based on earlier work by Siegel and Weil on modular forms Siegel1935 ; Siegel1951 ; Weil1964 ; Weil1965 and Segal and Shale on the quantization of the boson field Segal1959 ; Segal1961 ; Shale1962 .
@@
Here $\pi_{\lambda}$ are representations of $S_{g}$, and $\rho_{\lambda}$ are representations of $GL(n,{\mathbb{C}})$. They are both labeled by Young tableaux with $g$ boxes. For example, in the case of just two factors, we have $V\otimes V=S^{2}V+\Lambda^{2}V$, where $S^{2}V$ is the symmetric part (upon which $S_{2}$ acts trivially), and $\Lambda^{2}V$ is the antisymmetric part (upon which $S_{2}$ acts by a sign). Here $\pi_{\lambda}$ are the two one-dimensional irreps of $S_{2}$ and $\rho_{\lambda}$ are $n(n\pm 1)/2$ dimensional irreducible representations of $GL(n,{\mathbb{C}})$. [^15] Schur–Weyl duality has no direct relation to averages discussed in this work. However, see Maxfield:2023mdj for a setting where Schur–Weyl duality is central to characterizing the Hilbert space in a toy model of quantum gravity.
@@ 3.1 Generalities
Howe duality applies when $\mathcal{S}$ acts on a space $W$ and when its subgroups $G$ and $H$ form a dual pair, which requires that they be 1) reductive groups and 2) mutually centralizing inside $\mathcal{S}$. A group $G$ acts reductively on $W$ if it breaks $W$ into a finite sum of $G$-invariant subspaces. The centralizer of a subgroup $G\subset\mathcal{S}$ is the set of elements in $\mathcal{S}$ which commute with all $g\in G$. Mutually centralizing means that $G$ is the centralizer of $H$ and vice versa.
Howe gave a classification of reductive dual pairs Howe1979 ; Howe:1987tv ; Howe1989a :
@@
where $\mathbf{e}_{i}$ is a $2n\times g$ matrix with zeros except a single $1$ at a position determined by $1\leq i\leq 2gn$ (say row by row), and $\hat{X}_{i}$ is the operator $\hat{X}$ of (61) acting on the $i^{\text{th}}$ qudit. In the case of $F={\mathbb{F}}_{p}$, each “creation operator” can only act $p$ times, and $\omega$ has dimension $p^{2ng}$. The oscillator representation $\omega$ is then isomorphic as a vector space to $\mathcal{X}$ introduced above.
The oscillator representation $\omega$ is an irreducible representation of the Heisenberg–Weyl group $H_{2ng}$. It also transforms projectively under the group of (outer) automorphisms of $H_{2ng}$, which is $Sp(4ng,F)$ – these are the canonical transformations of quantum mechanics. In the case $F={\mathbb{R}}$, $\omega$ is a projective representation of $Sp(4ng,F)$ but a linear representation of the double cover $\widetilde{Sp}$ of the symplectic group, which is called the metaplectic group, and the pair $G\times H$ lifts to a pair of subgroups $\tilde{G}\times\tilde{H}$ of $\widetilde{Sp}$. [^16] For finite fields, it is possible to write down a genuine (non-projective) representation by including suitable phases, but this is impossible for $\mathbb{R}$}. For the field $\mathbb{C}$}, the oscillator representation is a genuine representation of $Sp(4gn,{\mathbb{C}})$ Howe1979 . A review of the details of this construction can be found in appendix A, where we give an introduction to the oscillator representation for both $\mathbb{R}$} and ${\mathbb{F}}_{p}$.
@@
[…]
@@ 3.2 Finite-field Howe duality
where $\Theta(\pi)$ is not necessarily an irrep, meaning that the correspondence between irreps is not one-to-one. [^19] Some work has gone into understanding if there is a natural choice for a canonical representative irrep in $\Theta(\pi)$ which makes the correspondence one-to-one, see Gurevich2016 ; Aubert2016 ; Pan2020 . Since all involved spaces are finite, it is always possible to write the decomposition as (110) (if one allows for $\Theta(\pi_{i})=0$ for some $\pi_{i}$). The non-trivial part is therefore to characterize the map $\Theta$. A conjecture about $\Theta$ for a subclass of irreps, so-called unipotent representations, that includes the singlet (trivial representation) was given in Aubert1996 and proven in Pan2019 ; Ma2022 . We give more details on this below, but first we formulate our main argument.
@@ 3.2.1 Proof of equality “Ensemble average equals Poincaré series”
[…]
It is easy to see that $\phi_{\text{A}}$ and $\phi_{\text{P}}$ are both invariant under both $H$ and GG. [^20] The invariance of $\phi_{\text{A}}$ under $G$ and $\phi_{\text{P}}$ under $H$ are immediate since they are group averages. The invariance of $\phi_{\text{A}}$ under $H$ follows from modular invariance of each $\psi_{C}$. Finally, the invariance of $\phi_{\text{P}}$ under $G$ follows from commutativity of actions of $G$ and $H$ and the fact that $\psi_{0}$ is invariant under $G$. They are also both non-zero. [^21] Neither $\phi_{\text{A}}$ nor $\phi_{\text{P}}$ can vanish. Each term contributing to $\phi_{\text{A}}$ has a positive scalar product with $\psi_{0}$; each term contributing to $\phi_{\text{P}}$ has a positive scalar product with $\psi_{\mathcal{C}}$ for any code $\mathcal{C}$. Then we invoke the following lemma, proven below.
@@ 3.2.2 Theta map and proof of the lemma
We now introduce some definitions following Pan2020 , which will allow us to explain (114) from the statements of Aubert1996 ; Pan2019 ; Ma2022 . An important subclass of representations of both orthogonal and symplectic groups over ${\mathbb{F}}_{p}$ ($p$ odd) are so-called unipotent representations [^22] The definition of unipotent is complicated, see Definition 7.8 of Deligne1976 ; for us it suffices to know that this is a subclass of representations which includes the singlet of the respective groups. and can be labeled by “symbols.” This follows older results by Lusztig Lusztig1977 . A symbol is a pair of collections of strictly decreasing numbers (or empty set),
@@
[…]
@@
Unipotent representations map to unipotent representations under the $\Theta$ map. To give an explicit description of how, we change labels from symbols to bipartitions. A bipartition is a pair of Young tableaux, equivalently a pair of integer collections $\kappa_{1},\ldots,\kappa_{m}]$, $\kappa_{i}\geqslant\kappa_{i+1}$. Restricting to the case of $\operatorname{def}\Lambda=1$ for $Sp$ and $\operatorname{def}\Lambda=0$ for $O^{+}$, we can map symbols to bipartitions by
@@
Let us now apply this consideration to the singlet (trivial) irrep. In terms of bi-partitions, the trivial irrep of $O^{+}(2n,{\mathbb{F}}_{p})$ has $(\kappa,\lambda)=([n],\emptyset)$, where $[n]$ denotes the Young tableau with a single row with $n$ boxes – see proposition 2.9 of Pan2020 . By checking the conditions in (121), we then find
@@ 3.3 Explicit example: $p=3$, $n=1$, $g=1$
Having discussed the general theory, we are ready to give an example with all the details spelled out. We will consider the simplest case, $p=3$, length $n=1$ and genus $g=1$. In appendix E, we give four additional examples: increasing each of the parameters (thus [png]=[511],[321],[312][png]=[511],\,[321],\,[312]), and finally the case $[png]=[322]$, where the sum (122) contains three terms instead of two as in the other examples.
@@
$W_{\mathcal{C}_{1}}=\psi_{00}+\psi_{10}+\psi_{20}\,,\qquad W_{\mathcal{C}_{2}}=\psi_{00}+\psi_{01}+\psi_{02}\,$$\qquad$ (124)
@@
$SL_{2,3}$ $[1]$ $[4]$ $4]^{\prime} [$$1]^{\prime} [$$^{\prime\prime} [4]′′′[4]^{\prime\prime\prime}$ $[6]$
$r_{1}$ 11 11 11 11 11 11 11
$r_{1}^{\prime}$ 11 $\zeta^{2}$ $\zeta$ 11 $\zeta$ $\zeta^{2}$ 11
$\bar{r}_{1}^{\prime}$ 11 $\zeta$ $\zeta^{2}$ 11 $\zeta^{2}$ $\zeta$ 11
$r_{2}$ 22 11 11 2-2 1-1 1-1 0
$r_{2}^{\prime}$ 22 $\zeta$ $\zeta^{2}$ 2-2 $\zeta^{2}$ $\zeta$ 0
$\bar{r}_{2}^{\prime}$ 22 $\zeta^{2}$ $\zeta$ 2-2 $\zeta$ $\zeta^{2}$ 0
$r_{3}$ 33 0 0 33 0 0 1-1
$O_{2,3}^{+}$ II $\left(\begin{smallmatrix}0&2\\ 2&0\end{smallmatrix}\right)$ $\left(\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right)$ $2I$
$\pi_{1}$ 11 11 11 11
$\pi_{1}^{\prime}$ 11 1-1 1-1 11
$\pi_{1}^{\prime\prime}$ 11 11 1-1 1-1
$\pi_{1}^{\prime\prime\prime}$ 11 1-1 11 1-1
@@
The larger symplectic group that acts on $\mathcal{X}$ is $\mathrm{Sp}(4,{\mathbb{F}}_{3})$. It has 51 84051\,840 elements and 34 irreps. Four of these are “minimal” (see appendix B) of dimension $4,4,5,5$, the other non-trivial irreps have dimensions between $6$ and $81$. Our space $\mathcal{X}$, which is isomorphic to the oscillator representation, is here a direct sum of one irrep with dimension $5$ (charge-conjugation even) and one irrep with dimension $4$ (charge-conjugation odd) under $Sp(4,{\mathbb{F}}_{3})$.
@@
We finish this section with a comment on modular invariants. We see that the following two spaces,
$\pi_{1}\otimes r_{1}$ $\leftrightarrow\operatorname{span}\{2\psi_{00}+\psi_{01}+\psi_{02}+\psi_{10}+\psi_{20}\},$ (134)
$\pi_{1}^{\prime}\otimes r_{1}$ $\leftrightarrow\operatorname{span}\{\psi_{01}+\psi_{02}-\psi_{10}-\psi_{20}\},$ (135)
@@
This idea that the enumerator polynomials of codes span all invariants of the MacWilliams identities (generally denoted Clifford groups) is an important concept in the study of error-correcting codes Nebe2006book . However, it is known that codes cannot be used to generate all modular forms (at large enough genus), even in the chiral case, as discussed in SalvatiManni1986 ; Runge1996 ; Oura2008 .